# Path-based Explanation for Knowledge Graph Completion

### v1

@@ Abstract.
Graph Neural Networks (GNNs) have achieved great success in Knowledge Graph Completion (KGC) by modelling how entities and relations interact in recent years. However, the explanation of the predicted facts has not caught the necessary attention. Proper explanations for the results of GNN-based KGC models increase model transparency and help researchers develop more reliable models. Existing practices for explaining KGC tasks rely on instance/subgraph-based approaches, while in some scenarios, paths can provide more user-friendly and interpretable explanations. Nonetheless, the methods for generating path-based explanations for KGs have not been well-explored. To address this gap, we propose Power-Link, the first path-based KGC explainer that explores GNN-based models. We design a novel simplified graph-powering technique, which enables the generation of path-based explanations with a fully parallelisable and memory-efficient training scheme. We further introduce three new metrics for quantitative evaluation of the explanations, together with a qualitative human evaluation. Extensive experiments demonstrate that Power-Link outperforms the SOTA baselines in interpretability, efficiency, and scalability.
@@
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Reasoning about belief and knowledge
@@ 1. Introduction
Providing explanations for GNN-based KGC models is a novel and challenging task. Given the nature of KGs in modeling structured real-world information, their end users often have little or no background in AI or ML. Therefore, the explanations need to be more aligned with human intuitions. Existing approaches attempt to explain KGs from triplet or subgraph perspectives, but the problem of what constitutes a good explanation of KGs is more complex than on homogeneous graphs, especially in terms of satisfying the “stakeholders’ desiderata” (Langer et al., 2021). In contrast to the widely adopted instance/subgraph-based explanations, we focus on generating the explanations of GNN-based KGC models from the perspective of paths in this work:
@@
Figure 1 illustrates three challenges of applying instance-based or subgraph-based GNN explanations from homogeneous graphs to KGs: (1) Interpretability: For KGs, explanations should reveal the connections between a pair of entities given a specific relation through paths on KGs. However, previous approaches produce disconnected subgraphs that do not show this connected pattern (e.g., left from Figure 1). These subgraphs are not suitable for human interpretation and exploration. (2) Sufficiency: Paths contain sufficient information to explain whether a given triplet is factual. From Figure 1, for explaining triplet $\langle New\_York,located\_in,North\_American\rangle$, the subgraph-based explanation brings in a noisy and irrelevant fact $\langle New\_York,has\_team,Knicks\rangle$. The instance-based one only extracts a related fact $\langle New\_York,most\_popular\_city,United\_States\rangle$ but loses sufficient information for users to interpret. Therefore, neither method is suitable for KGs where concise and informative semantics are preferred. (3) Scalability: Compared to general subgraphs, paths form a much smaller search space for optimal subgraph explanations since paths are simpler and scale linearly in the number of graph edges. Therefore, paths can better fit with KGs that usually come with massive entities. To our best knowledge, there is only one approach, PaGE-Link (Zhang et al., 2023) that explores path-based solutions to the GNN explanation problem. But PaGE-Link focuses on heterogeneous graphs without considering the relation types and the scalability of the method, which makes it unsuitable to be directly applied to KGs. Therefore, the question is still underexplored on KGs.
@@
Given the important roles that paths play in the explanation of KGC tasks, we propose Power-Link to tackle the aforementioned challenges, which aims to enhance the interpretability, sufficiency, and scalability of the explanations by providing explainable paths for the GNN-based KGC models. To the best of our knowledge, this is the first work specifically designed for explaining KGC tasks via paths. Figure 2 depicts the overall framework of Power-Link. 1) We first extract a k-hop subgraph around the source and target nodes and prune it to eliminate noisy neighbours. 2) A Triplet Edge Scorer (TES) is proposed to measure the importance of each edge along the candidate paths by leveraging the node and edge type information, which is specifically designed for the KG scenario. 3) PaGE-Link employs the shortest-path searching algorithm to emphasize paths in each iteration, which is computationally intensive and not applicable to large-scale KGs. It also introduces the accumulation of on-path errors during the learning process. To alleviate this issue, we propose to replace the shortest-path searching algorithm with a specially-designed graph-Power-based method to amplify explainable on-path edges. Our method is highly parallelisable with GPUs and consumes minimal memory resources, enabling easy scaling to large KGs. The graph power-based approach also provides an interface for end users to select the desired path length of each explanation manually. 4) By combining the path-amplifying loss and the mutual-information-based loss, Power-Link generates interpretable paths to explain the prediction of GNN-based KGC models in an efficient manner. 5) To better measure the performance of explanation methods on KGs from a quantitive perspective, we also propose three proper metrics, $Fidelity+$, $Fidelity-$, and $H\Delta R$, that reflect how good an explanation is in the empirical evaluation.
@@ 2.3. Paths on Graphs
We briefly review the methods which leverage paths on graphs as crucial parts of their algorithms. Paths are often used to represent the structure and semantics of graphs, such as Katz index (Katz, 1953), graph distance (Bunke and Shearer, 1998), SimRank (Jeh and Widom, 2002), and PathSim (Sun et al., 2011). They can also reveal the underlying relationship between a pair of nodes, and therefore bring more explanability to the methods. (Luo et al., 2021) propose a context path-based GNN that recursively embeds the paths connecting nodes into the node embedding with attention mechanisms. In the same field, Zhu et al. (2021) utilizes the idea from the neural bellman-ford algorithm to construct a general path-based framework NBFNet for both link prediction and KGC. Zhu et al. (2022) further enhance the scalability of NBFNet by proposing to use A*** algorithm for path constructions. Chang et al. (2023) propose to augment the knowledge used in the KGC task from the counterfactual view and enhance the explainability of the completion results. While these methods can generate non-trivial path-based explanations as a side-product, they cannot be applied to the black-box explanation setting in explainable artificial intelligence (XAI), where the pre-trained KG embeddings are given. Therefore, we argue that paths are still valuable for providing interpretable explanations for KGC under the local post-hoc scenario.
@@ 4. Problem Formulation
We focus on the task of generating path-based explanations for a predicted fact between a pair of entities $e_{i}$ and $e_{j}$. Given a trained GNN-based KGC model Φ(,)normal-$\Phi(\cdot,\cdot)$, a computation graph ${\mathcal{G}}_{c}$ of $\hat{h},\hat{t}$ that extracted from the target KG ${\mathcal{G}}$, we find an explanation ${\mathcal{P}}=$ $\pi|\pi$ is a $\hat{h}-\hat{t}$ path with maximum length $L$. By proper construction of optimization on candidates from $\pi\in{\mathcal{P}}$, we aim to generate concise and influential paths as explanations for the prediction.
@@ 5.2. Path-Enforcing Learning
We first briefly discuss the problem we find in the learning process of PaGE-Link. They optimize an explanatory weighted mask and enhance the path-forming explanations by simultaneously forcing the on-path weights to increase and the off-path weights to decrease. The potential paths are selected using the shortest-path searching algorithm, whose cost matrix is designed based on the weighted mask and restrictions on the node degree. First, the shortest-path searching algorithm is not parallelisable, which is slow when scaled to large KGs. Second, we argue that this optimizing strategy may introduce noise at the early training stage. In the beginning, higher weights in the mask can be assigned to trivial edges because of incomplete training, while meaningful edges are ignored. Therefore, when applying shortest-path searching on the underfitting weighted mask, the algorithm may strengthen the meaningless paths and weaken the important explanatory paths. This introduces noise that can be propagated through epochs. We also observe that PaGE-Link often generates similar explanations as the GNNExplainer, which is not equipped with a path-enforcing training strategy. We consider this result from the early perturbations in the training process, which hinders the algorithm from finding explanatory paths.
@@ 5.3. Path Generation
Algorithm 1 describes the full process of Power-Link. The learning process is highly parallelizable and can be accelerated with GPUs. Since the score matrix ${\mathcal{M}}$ and the adjacency matrix are usually sparse for KGs, by utilizing a sparse matrix, we are able to enjoy even faster and memory-efficient computation. We also provide a complexity analysis of Power-Link and other methods in the Experiments section.
@@
9:    for $l<L-1$ do
10:         Compute the power of the power vector $\bm{u}^{(l)}=\bm{u}^{(l-1)}\ {\mathcal{M}}$;
11:         Compute the average edge probability by Eq.(8);
12:         Compute the on-path probability $\bm{u}^{(l)}_{\hat{k}}$;
13:         Record the on-path probability $P_{on}=P_{on}+\bm{u}^{(l)}_{\hat{k}}$
14:    end for
15:    Compute the total loss by Eq.(12);
16:    Update $TES$ through backpropagation;
17:end while
18:return ${\mathcal{P}}\leftarrow Dijkstra(\hat{h},\hat{t},{\mathcal{G}}_{c}^{k},{\mathcal{M}})$.
@@ 6.2. Evaluation Metrics
Sparsity. As the Fidelity score is often positively correlated with Sparsity, we consider Sparsity as one of our evaluation metrics. We measure the soft sparsity of the explanatory mask in our experiment. For a good explanation, the explanatory mask should be highly condensed on the meaningful edges, therefore rendering high sparsity.
@@ 6.4. Ablation Studies
Comparison with PaGE-Link on heterogeneous graphs. Since the path-enforcing learning module in Power-Link can also be applied to heterogeneous graphs, for a better comparison, we reproduce the experiments in PaGE-Link with our proposed learning module. In both AugCitation and UserItemAttr datasets, Power-Link achieves an improvement of 0.05 on the AUC score and is around 1.5 times faster. The result aligns with previous experiments, showing that Power-Link is more precise and efficient. The specific statistics can be found in the Appendix.
@@ 7. Conclusion
Appendix: Path-based Explanation for Knowledge Graph Completion
@@ Appendix C Statistics of datasets
Appendix D Reproduce PaGE-Link experiments with Power-Link

### v2

@@ Abstract.
Graph Neural Networks (GNNs) have achieved great success in Knowledge Graph Completion (KGC) by modelling how entities and relations interact in recent years. However, the explanation of the predicted facts has not caught the necessary attention. Proper explanations for the results of GNN-based KGC models increase model transparency and help researchers develop more reliable models. Existing practices for explaining KGC tasks rely on instance/subgraph-based approaches, while in some scenarios, paths can provide more user-friendly and interpretable explanations. Nonetheless, the methods for generating path-based explanations for KGs have not been well-explored. To address this gap, we propose Power-Link, the first path-based KGC explainer that explores GNN-based models. We design a novel simplified graph-powering technique, which enables the generation of path-based explanations with a fully parallelisable and memory-efficient training scheme. We further introduce three new metrics for quantitative evaluation of the explanations, together with a qualitative human evaluation. Extensive experiments demonstrate that Power-Link outperforms the SOTA baselines in interpretability, efficiency, and scalability. The code is available at https://github.com/OUTHIM/power-link
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671683isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Reasoning about belief and knowledge
@@ 1. Introduction
[…]
Example of the advantage that path-based explanation has over subgraph-based ones.
@@
[…]
The overall framework of the proposed Power-Link. Given a KG and a trained KGC model as inputs, we aim to generate interpretable paths to explain why the KGC model predicts a target triplet is factual.
Providing explanations for GNN-based KGC models is a novel and challenging task. Given the nature of KGs in modeling structured real-world information, their end users often have little or no background in AI or ML. Therefore, the explanations need to be more aligned with human intuitions. Existing approaches attempt to explain KGs from triplet or subgraph perspectives. Still, the problem of what constitutes a good explanation of KGs is more complex than on homogeneous graphs, especially in terms of satisfying the “stakeholders’ desiderata” (Langer et al., 2021). In contrast to the widely adopted instance/subgraph-based explanations, we focus on generating the explanations of GNN-based KGC models from the perspective of paths in this work:
@@
Figure 1 illustrates three challenges of applying instance-based or subgraph-based GNN explanations from homogeneous graphs to KGs: (1) Interpretability: For KGs, explanations should reveal the connections between a pair of entities given a specific relation through paths on KGs. However, previous approaches produce disconnected subgraphs that do not show this connected pattern (e.g., left from Figure 1). These subgraphs are not suitable for human interpretation and exploration. (2) Sufficiency: Paths contain sufficient information to explain whether a given triplet is factual. From Figure 1, for explaining triplet $\langle New\_York,located\_in,North\_American\rangle$, the subgraph-based explanation brings in a noisy and irrelevant fact $\langle New\_York,has\_team,Knicks\rangle$. The instance-based one only extracts a related fact $\langle New\_York,most\_popular\_city,United\_States\rangle$ but loses sufficient information for users to interpret. Therefore, neither method is suitable for KGs where concise and informative semantics are preferred. (3) Scalability: Compared to general subgraphs, paths form a much smaller search space for optimal subgraph explanations since paths are simpler and scale linearly in the number of graph edges. Therefore, paths can better fit with KGs that usually come with massive entities. To our best knowledge, there is only one approach, PaGE-Link (Zhang et al., 2023) that explores path-based solutions to the GNN explanation problem. But PaGE-Link focuses on heterogeneous graphs without considering the relation types and the scalability of the method, which makes it unsuitable to be directly applied to KGs. Therefore, the question of KGs is still underexplored.
@@
Given the important roles that paths play in the explanation of KGC tasks, we propose Power-Link to tackle the aforementioned challenges, which aims to enhance the interpretability, sufficiency, and scalability of the explanations by providing explainable paths for the GNN-based KGC models. To the best of our knowledge, this is the first work specifically designed to explain KGC tasks via paths. Figure 2 depicts the overall framework of Power-Link. 1) We first extract a k-hop subgraph around the source and target nodes and prune it to eliminate noisy neighbours. 2) A Triplet Edge Scorer (TES) is proposed to measure the importance of each edge along the candidate paths by leveraging the node and edge type information, which is specifically designed for the KG scenario. 3) PaGE-Link employs the shortest-path searching algorithm to emphasize paths in each iteration, which is computationally intensive and not applicable to large-scale KGs. It also introduces the accumulation of on-path(edges on the paths) errors during the learning process. To alleviate this issue, we propose to replace the shortest-path searching algorithm with a specially-designed graph-Power-based method to amplify explainable on-path edges. Our method is highly parallelisable with GPUs and consumes minimal memory resources, enabling easy scaling to large KGs. The graph power-based approach also provides an interface for end users to select the desired path length of each explanation manually. Power-Link generates interpretable paths to explain the prediction of GNN-based KGC models in an efficient manner. 5) To better measure the performance of explanation methods on KGs from a quantitive perspective, we also propose three proper metrics, $Fidelity+$, $Fidelity-$, and $H\Delta R$, that reflect how good an explanation is in the empirical evaluation.
@@ 2.3. Paths on Graphs
We briefly review the methods which leverage paths on graphs as crucial parts of their algorithms. Paths are often used to represent the structure and semantics of graphs, such as Katz index (Katz, 1953), graph distance (Bunke and Shearer, 1998), SimRank (Jeh and Widom, 2002), and PathSim (Sun et al., 2011). They can also reveal the underlying relationship between a pair of nodes, and therefore bring more explanability to the methods. (Luo et al., 2021) propose a context path-based GNN that recursively embeds the paths connecting nodes into the node embedding with attention mechanisms. In the same field, Zhu et al. (2021) utilizes the idea from the neural bellman-ford algorithm to construct a general path-based framework NBFNet for both link prediction and KGC. Zhu et al. (2022) further enhance the scalability of NBFNet by proposing to use A* algorithm for path constructions. Chang et al. (2023) propose to augment the knowledge used in the KGC task from the counterfactual view and enhance the explainability of the completion results. While these methods can generate non-trivial path-based explanations as a side-product, they cannot be applied to the black-box explanation setting in explainable artificial intelligence (XAI), where the pre-trained KG embeddings are given. Therefore, we argue that paths are still valuable for providing interpretable explanations for KGC under the local post-hoc scenario.
@@ 4. Problem Formulation
We focus on the task of generating path-based explanations for a predicted fact between a pair of entities $e_{i}$ and $e_{j}$. Given a trained GNN-based KGC model $\Phi(\cdot,\cdot)$, a computation graph ${\mathcal{G}}_{c}$ of $\hat{h},\hat{t}$ that extracted from the target KG ${\mathcal{G}}$, we find an explanation ${\mathcal{P}}=$ $\pi|\pi$ is a $\hat{h}-\hat{t}$ path with maximum length $L$. By proper construction of optimization on candidates from $\pi\in{\mathcal{P}}$, we aim to generate concise and influential paths as explanations for the prediction.
@@ 5.2. Path-Enforcing Learning
We first briefly discuss the problem we find in the learning process of PaGE-Link. They optimize an explanatory weighted mask and enhance the path-forming explanations by simultaneously forcing the on-path (edges on the paths between the target entities) weights to increase and the off-path (edges not on the paths between the target entities) weights to decrease. The potential paths are selected using the shortest-path searching algorithm, whose cost matrix is designed based on the weighted mask and restrictions on the node degree. First, the shortest-path searching algorithm is not parallelisable, which is slow when scaled to large KGs. Second, we argue that this optimizing strategy may introduce noise at the early training stage. In the beginning, higher weights in the mask can be assigned to trivial edges because of incomplete training, while meaningful edges are ignored. Therefore, when applying shortest-path searching on the underfitting weighted mask, the algorithm may strengthen the meaningless paths and weaken the important explanatory paths. This introduces noise that can be propagated through epochs. We also observe that PaGE-Link often generates similar explanations as the GNNExplainer, which is not equipped with a path-enforcing training strategy. We consider this result from the early perturbations in the training process, which hinders the algorithm from finding explanatory paths.
@@ 5.3. Path Generation
Algorithm 1 describes the full process of Power-Link. The learning process is highly parallelizable and can be accelerated with GPUs. Since the score matrix ${\mathcal{M}}$ and the adjacency matrix are usually sparse for KGs, by utilizing the sparse matrix operations, we are able to enjoy even faster and memory-efficient computation. We also provide a complexity analysis of Power-Link and other methods in the Experiments section.
@@
9:    Obtain the position $(\hat{i},\hat{k})$ of the target triplet in ${\mathcal{M}}$
10:    Initialize the power vector $\bm{u}={\mathcal{M}}_{\hat{i}:}$
11:    for $l<L-1$ do
12:         Compute the power of the power vector $\bm{u}^{(l)}=\bm{u}^{(l-1)}\ {\mathcal{M}}$;
13:         Compute the average edge probability by Eq.(8);
14:         Accumulate the target on-path probability $P_{on}=P_{on}+\bm{u}^{(l)}_{\hat{k}}$
15:    end for
16:    Average the on-path probability $P_{on}=\frac{P_{on}}{L-1}$
17:    Compute the total loss by Eq.(12);
18:    Update $TES$ through backpropagation;
19:end while
20:return ${\mathcal{P}}\leftarrow Dijkstra(\hat{h},\hat{t},{\mathcal{G}}_{c}^{k},{\mathcal{M}})$.
@@ 6.1. Experimental Setup
[…]
The explanations (green and blue arrows) by different explainers for the prediction $\langle Limitless,release\_in,Argentina\rangle$ (dashed red). Power-Link explains the fact by the $release\_in$ relationship, whereas baseline explanations are less interpretable.
@@ 6.2. Evaluation Metrics
[…]
The GPU memory usage of Power-Link during the explaining process against the number of edges in each graph. 500 samples are explained. For better illustration, we only record the memory usage of the explaining module. The memory occupied by the KGC model is ignored.
Sparsity. As the Fidelity score is often positively correlated with Sparsity, we consider Sparsity as one of our evaluation metrics. In our experiment, we measured the soft sparsity of the explanatory mask. For a good explanation, the explanatory mask should be highly condensed on the meaningful edges, therefore rendering high sparsity.
@@ 7.1. KGC Explanation Results
Table 4. Ablation study on the path-enforcing module. (-MI) indicates the model is explained with the mutual information loss only. The asterisk * indicates the results from Table 2.
[…]
RGCN-DistMult-MI 0.356 0.168 0.834 0.173 0.308 0.396
WGCN-DistMult-MI 0.325 0.397 0.640 0.599 0.683 0.739
CompGCN-DistMult-MI 0.0411 0.0299 0.6556 0.072 0.126 0.194
RGCN-DistMult 0.377 0.186 0.784 0.234 0.354 0.432
WGCN-DistMult 0.646 0.294 0.672 0.629 0.772 0.836
CompGCN-DistMult 0.043 0.025 0.506 0.092 0.170 0.220
@@ 7.2. Ablation Studies
We conduct ablation studies using the KGC models trained on the FB15K-237 dataset to verify the impact of the design components in our proposed method.
The path-enforcing module. As shown in Table 4, we observed an obvious drop in path explanation performance when the path-enforcing module was removed. This is as expected since there is no strengthening on the impact of paths when explanations are generated without the path-enforcing. We expect the gap to be larger for denser graphs.
The combination method in TES. Table 5 shows that using the Euclidean combination method yields better explanation performance than Concatenation on RGCN-based KGC models. The superiority is more obvious for RGCN-TransE, which uses the energy-based loss function during KGC training. On the contrary, Euclidean downgrades the performance on WGCN-based KGC models.
The mutual information loss. We test the impact of mutual information loss(MIL) in explaining the WGCN-Distmult model. Table 6 shows that without the mutual information loss, there is a significant drop in the explanation performance. The MIL is the key component that enables the explanation method to adjust the weighted mask of the KG according to the prediction of the KGC model. Without MIL, the path-enforcing module only strengthens the edges on the paths between the head and tail nodes without considering their impacts on the model’s prediction.
The order of powering. Table 7 presents the explanation performance of Power-Link when we change the order of powering in the path-enforcing module. Interestingly, we find that increasing the powering order brings better performance and vice versa. We believe that increasing the powering order expands the range of length of the paths that the module strengthens. This can enable the generation of more explanatory paths that contain more important nodes. Even though, it is worth noting that longer paths may be less meaningful to the users, and higher powering order requires more computational power.
Comparison with PaGE-Link on heterogeneous graphs. Since the path-enforcing learning module in Power-Link can also be applied to heterogeneous graphs, for a better comparison, we reproduce the experiments in PaGE-Link with our proposed learning module. In both AugCitation and UserItemAttr datasets, Power-Link achieves an improvement of 0.05 on the AUC score and is around 1.5 times faster. The result aligns with previous experiments, showing that Power-Link is more precise and efficient. The specific statistics can be found in the Appendix Table 12.
Table 5. Ablation study on the combination methods in TES. The asterisk * indicates the results from Table 2.
[…]
RGCN-TransE-euc 0.711 0.204 0.906 0.480 0.640 0.717
RGCN-DistMult-euc 0.266 0.133 0.763 0.334 0.455 0.468
RGCN-ConvE-euc 0.189 0.206 0.936 0.444 0.542 0.590
WGCN-TransE-euc 0.761 0.444 0.648 0.382 0.538 0.656
WGCN-DistMult-euc 0.628 0.375 0.580 0.594 0.737 0.822
WGCN-ConvE-euc 0.794 0.377 0.581 0.416 0.650 0.726
RGCN + TransE-concat 0.699 0.088 0.531 0.074 0.174 0.234
RGCN + DistMult-concat 0.377 0.186 0.784 0.234 0.354 0.432
RGCN + ConvE-concat 0.435 0.111 0.832 0.190 0.326 0.396
WGCN + TransE-concat 0.775 0.150 0.753 0.376 0.582 0.648
WGCN + DistMult-concat 0.646 0.294 0.672 0.629 0.772 0.836
WGCN + ConvE-concat 0.416 0.287 0.720 0.474 0.646 0.762
Table 6. Ablation study on the mutual information loss. (-PS) indicates that only path-searching loss is used during the explanation(without mutual information loss). The asterisk * indicates the results from Table 2.
[…]
WGCN-DistMult-PS 0.013 0.021 0.938 0.168 0.220 0.238
WGCN-DistMult 0.646 0.294 0.672 0.629 0.772 0.836
Table 7. Ablation study on the powering order. The asterisk * indicates the results from Table 2.
[…]
RGCN-DistMult 4 0.367 0.172 0.803 0.224 0.376 0.440
RGCN-DistMult 2 0.172 0.175 0.954 0.200 0.296 0.350
RGCN-DistMult 3 0.377 0.186 0.784 0.234 0.354 0.432
7.3. Visualization and Performance Analysis
@@ 8. Conclusion
Acknowledgements.
This work was partially supported by NSFC Grant No. 62206067 and Guangzhou-HKUST(GZ) Joint Funding Scheme 2023A03J0673.
@@ Appendix C Statistics of datasets
Appendix D Metrics of the KGC models
We provide the metrics of the KGC models used for explanation in Table 11.
Table 11. Metrics of the KGC models used for explanation.
Dataset Model MRR Hit@1 Hit@3 Hit@10
FB15K-237 CompGCN-ConvE 0.352 0.261 0.385 0.536
CompGCN-DistMult 0.343 0.252 0.376 0.524
CompGCN-TransE 0.230 0.246 0.368 0.510
RGCN-ConvE 0.430 0.376 0.461 0.524
RGCN-DistMult 0.380 0.327 0.414 0.471
RGCN-TransE 0.210 0.132 0.228 0.474
WGCN-ConvE 0.319 0.235 0.349 0.487
WGCN-DistMult 0.272 0.200 0.295 0.415
WGCN-TransE 0.288 0.213 0.315 0.434
WN18RR CompGCN-ConvE 0.431 0.395 0.439 0.515
CompGCN-DistMult 0.431 0.395 0.439 0.515
CompGCN-TransE 0.206 0.126 0.276 0.507
RGCN-ConvE 0.464 0.389 0.472 0.533
RGCN-DistMult 0.331 0.311 0.348 0.402
RGCN-TransE 0.182 0.132 0.203 0.396
WGCN-ConvE 0.449 0.413 0.467 0.521
WGCN-DistMult 0.320 0.283 0.337 0.383
WGCN-TransE 0.118 0.032 0.165 0.284
Appendix E Reproduce PaGE-Link experiments with Power-Link